OpenStudy (anonymous):

? Solve for x. -3 + 8x - 5 = -8

OpenStudy (anonymous):

x = 0

OpenStudy (anonymous):

-3+8x-5=-8 8x=-8+5+3 8x=0 x=0/8

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

-3 + 8x -5 = -8 first, put the -3 on the right side: 8x -5 = -8 + 3 then, put the -5 on the right side too: 8x = -8 + 3 + 5 simplify! once you solve for 8x, divide by 8 to get x!